OpenStudy (anonymous):

A store mixes Brazilian coffee worth $10 per kilogram and Turkish coffee worth $12 per kilogram. The mixture is to sell for $11 per kilogram. Find how much of each should be used to make a 468 kilogram mixture

OpenStudy (eyust707):

10*(# of Brazillian Kilos) + 12 * (# of Turkish kilos) = Cost Lets call the "# of Brazilian kilos" the letter "B" and "# of turkish kilos" the letter "T" So: 10*B + 12*T = Total Cost and Total Cost = 11 * 468 Also B + T = 468 Solve for B and T

OpenStudy (anonymous):

10B+12T=11*468 10B+12(468-B)=11*468 10B+12*468-12B=11*468 12*468-11*468=12B-10B B=234

OpenStudy (eyust707):

For this one however, because they chose a price that is exactly halfway between the two prices, the proportions of the two coffees will be 50/50. So half the mixture will be Brazilain and half will be Turkish.

OpenStudy (eyust707):

Which @sauravshakya shows =)

OpenStudy (anonymous):

thankyou all...I was wondering why i kept coming out with equal amounts...now i know why...lol

OpenStudy (anonymous):

@eyust707 I know B+T=468 But how did u simply get that?